Meir, Benson, and Lau are partners and share income and loss in a 2:3:5 ratio (in percents: Meir, 20%; Benson, 30%; and Lau, 50%). The partnership's capital balances are as follows: Meir, $88,000; Benson, $134,000; and Lau, $228,000. Benson decides to withdraw from the partnership.

Assume that Benson does not retire from the partnership described in Part 1. Instead, Rhode is admitted to the partnership on February 1 with a 25% equity. Prepare journal entries to record Rhode’s entry into the partnership under each separate assumption: Rhode invests (a) $150,000; (b) $109,500; and (c) $196,500. (Do not round your intermediate calculations.)

Record the admission of Rhode with an investment of $150,000 for a 25% interest in the equity.

Record the admission of Rhode with an investment of $109,500 for a 25% interest in the equity.

Record the admission of Rhode with an investment of $196,500 for a 25% interest in the equity.
Answer #1

Solution:

Journal Entries:

Transaction General Journal Debit Credit
(a) Cash $          150,000
Rhode Capital $      150,000
(Being new partner admission)
(b) Cash $          109,500
Meir Capital (2/10) $               6,075
Benson Capital(3/10) $               9,112
Lau Capital (5/10) $            15,188
Rhode Capital $      139,875
(Being new partner admission)
(c) Cash $          196,500
Meir Capital (2/10) $           6,975
Benson Capital(3/10) $        10,462
Lau Capital (5/10) $        17,438
Rhode Capital $      161,625
(Being new partner admission)

Working:

a) Rhode with an investment of $150,000 for a 25% interest in the equity.

Particular Amount$
Capital Before admission $        450,000.00
Add: Rhode capital (a) $        150,000.00
Total Capital (b) $        600,000.00
Rhode Share of Capital 25% (c =b*25%) $        150,000.00
Bonus (a-c) $                          -  

b) Rhode with an investment of $109,500 for a 25% interest in the equity.

Particular Amount$
Capital Before admission $      450,000.00
Add: Rhode capital (a) $      109,500.00
Total Capital (b) $      559,500.00
Rhode Share of Capital 25% (c =b*25%) $      139,875.00
Bonus to Rhode by existing partners (2:3:5) (d=a-c) $        30,375.00
Meir (2/10)*d $          6,075.00
Benson(3/10)*d $          9,112.50
Lau(5/10)*d $        15,187.50
Total Bonus $        30,375.00

c) Rhode with an investment of $196,500 for a 25% interest in the equity.

Particular Amount$
Capital Before admission $      450,000.00
Add: Rhode capital (a) $      196,500.00
Total Capital (b) $      646,500.00
Rhode Share of Capital 25% (c =b*25%) $      161,625.00
Bonus to existing partners by rhode (2:3:5) (d=a-c) $        34,875.00
Meir (2/10)*d $          6,975.00
Benson(3/10)*d $        10,462.50
Lau(5/10)*d $        17,437.50
Total Bonus $        34,875.00

4) Final Answer in journal entries are round off to nearest whole dollar amount.